Suppose a monopolist faces a demand curve P = 60 – 3Q. Marginal cost is constant and equal to 6 (MC = 6). If a per-unit excise tax, t = $9 per unit is levied on consumers, how much will consumers pay for each unit of the good after the tax?
Answer #1

Ans.

MC after imposition of tax = 6+t

= 6+9

MC = 15

TR = ( 60 - 3Q) Q

MR = 60 - 6Q

Profit Maximisation at point

MR = MC

60 - 6Q = 15

6Q = 45

Q = 7.5

P = 60 - 3(7.5)

P = 60 - 22.5

P = 37.5